A cylinder has a radius of 3 cm and a height of 8 cm. What is the longest segment, in centimeters, that would fit inside the cylinder?
Answer by checkley77(12844)

What you have is a right triangle with a base=2*3=6cm & a height=8cm.
The longest segment would be the hypotenuse of this triangle thus.
a^2+b^2=c^2 [c being the hypotenuse)
6^2+8^2=c^2
36+64=c^2
100=c^2
c=sqrt100
c=10 ans. for the longest segment that would fit inside the cylinder.